What is the geometric series of 36+6+1.....S9?
Answer by stanbon(75887)

What is the geometric series of 36+6+1.....S9?
----
a(1) = 36
r = 6/36 = 1/6
-----
Not sure what you mean by S9, but I'll assume it you mean a(9).
-------
a(9) = a(1)*r^8
==
a(9) = 36*(1/6)^8 = 1/6^6 = 0.00002143...
Cheers,
Stan H.